# Leveraging Diverse Semantic-based Audio Pretrained Models for Singing Voice Conversion

### v2

@@ 4.2 Performance of Different Semantic-based Features (EQ1)
Figure 3: The complementary role of diverse semantic-based features in melody modeling.
It can be observed that on the one hand, the underlying knowledge of different semantic-based models can be distinct and diverse: (1) To model melody (F0CORR and F0RMSE), we find that Whisper and ContentVec behave better than WeNet. This is mainly because that the auxiliary tasks of the weak-supervised Whisper and the self-supervised ContentVec will be beneficial for modeling more prosody-related signals. (2) To model lyrics (CER), Whisper is the best compared to the other two. It is supposed that the multilingual ASR tasks and the large-scale pretraining data enable Whisper to model more valid and robust semantic information. Besides, ContentVec will provide better semantic signals than WeNet, although the two are both pretrained on only speech. It reveals that the self-supervised pretrained models could be more robust than the classic supervised ASR model. (3) To measure the speaker-independency (SIM), the speaker similarity results of the three are comparable and all over 0.9, which is hard to rank from human perception (see our demo page). It means that when using only semantic-based features for SVC, the three can all be considered speaker-agnostic.
@@ 4.3 Performance of the DSFF-SVC framework (EQ2)
To verify the generalization and robustness of the proposed DSFF-SVC framework, we conduct the both recording studio and in-the-wild setting tasks for the three base models. Except for the semantic-based features, here we also use the auxiliary prosody features (including F0 and energy), to improve the melody modeling and the expressiveness of the system. The objective and subjective evaluation results can be seen in Table 4. It reveals that: (1) Conducting SVC under in-the-wild setting is more difficult than recording studio setting. Facing to the more complex acoustic environment, only using the supervised WeNet produces a very poor performance. Integrating with the weak-supervised Whisper and the self-supervised ContentVec are usually beneficial, especially for Whisper. (2) For the generalization ability, we can observe that for all the three base models, the proposed idea of diversion semantic-based features fusion is effective in most cases under both recording studio and in-the-wild settings. (3) For the robustness, particularly, under the more challenging in-the-wild conversion setting, the improvement of integrating diverse semantic-based features is more obvious, especially for the objective CER and the subjective metrics.
@@
WeNet + Whisper 0.924 77.2 14.9% 0.881 0.848 183.8 18.7% 0.867
WeNet + Whisper + ContentVec 0.931 75.5 16.2% 0.883 0.857 186.7 23.3% 0.868
@@
WeNet + Whisper 0.945 144.4 17.8% 0.890 0.920 86.9 35.2% 0.869
WeNet + Whisper + ContentVec 0.946 112.9 17.7% 0.886 0.921 79.5 32.3% 0.870
@@
WeNet + Whisper 0.943 49.5 15.2% 0.884 0.921 73.6 21.1% 0.865
WeNet + Whisper + ContentVec 0.940 55.2 15.7% 0.884 0.919 79.9 23.3% 0.867
(a) Objective evaluation results.
Semantic-based Features Recording Studio Setting In-the-Wild Setting
Naturalness ($\uparrow$) Similarity ($\uparrow$) Naturalness ($\uparrow$) Similarity ($\uparrow$)
@@
WeNet + Whisper 4.02 ±$\pm 0.18$ 3.13 ±$\pm 0.17$ 3.70 ±$\pm 0.18$ 2.86 ±$\pm 0.23$
WeNet + Whisper + ContentVec 4.14 ±$\pm 0.19$ 3.25 ±$\pm 0.18$ 3.71 ±$\pm 0.18$ 2.82 ±$\pm 0.23$
(b) MOS evaluation results (with 95% confidence interval) of DiffWaveNetSVC. The full scores of Naturalness and Similarity are 5 and 4.
Table 4: Results of the proposed DSFF-SVC framework.
@@
RTX ($\uparrow$) RTF ($\downarrow$) F0CORR ($\uparrow$) F0RMSE ($\downarrow$) CER ($\downarrow$) SIM ($\uparrow$)
Cross attention 191.9 2.57 0.896 88.6 16.4% 0.871
ReTrans 415.0 0.86 0.936 55.5 15.8% 0.875
Table 5: The comparison between the proposed ReTrans and cross attention for features fusion.
Compared with Table 3 and 4, there is also many interesting observations about introducing the auxiliary prosody features (F0 and energy): (1) After injecting such prosody signals into SVC models, the intelligibility (CER) has been improved a lot. For example, using WeNet features on DiffWaveNetSVC, the CER is decreased from 38.2% to 15.8%. In our opinions, this is because that the prosody in singing voice are expressive and characteristic, which is hard to be modeled without the explicit assistant of prosody features like F0. Therefore, using only semantic-based features, some tone- and intonation-related signals will not be learned well, leading to the worse intelligibility. (2) Except for the intelligibility, it is also notable about the change of conversion speaker similarity (SIM). After introduce the explicit assistant of F0 features, the melody-related metrics have been improved a lot and the converted singing voices are not out of tune any more. In the meantime, however, the speaker similarity is also harmed. We suppose that such auxiliary prosody features own some speaker-specific signals, such as the personalized vibrato patterns within F0 features. Therefore, just copying the original prosody features from the source audio could not be the best choice for SVC, although it is a common way in the most SVC works.
To verify the necessity and efficiency of the proposed ReTrans strategy for feature alignment of mismatched time resolutions, we select the classic cross attention Vaswani et al. (2017) as a baseline. Specifically, under DSFF-SVC framework, we utilize only WeNet to extract semantic-based features and try different strategies to align its resolution to that of F0 and energy features. We conduct the experiments on DiffWaveNetSVC and follow Shen et al. (2024) to adopt the cross attention for features fusion within the diffusion model.
Except for the quality evaluation for the conversion results, we also measure the computational cost of the both. Assume the duration of the whole training corpus is $d_{train}$, the duration of the evaluation samples to be converted is $d_{infer}$, the time to train an epoch is $t_{train}$, and the time to infer the evaluation samples is $t_{infer}$. We define RTX as $d_{train}/t_{train}$, which can measure the training efficiency. We define RTF as $t_{infer}/d_{infer}$, which can measure the inference latency. The computational platform is a single NVIDIA Tesla V100.
The comparison results between ReTrans and cross attention are displayed in Table 5. We can see that when fusing the features of mismatched time resolutions for the SVC task, our proposed ReTrans is slightly better than cross attention in terms of conversion quality. However, as a non-learning method, ReTrans does not introduce any additional parameters and greatly accelerates both training efficiency (RTX) and inference speed (RTF) much.
@@ 5 Conclusion and Future Work
This paper investigates the underlying knowledge within the three representatives from supervised, weak-supervised, and self-supervised semantic-based pretrained models for singing voice conversion. We discover that the capabilities including modeling melody and lyrics of different semantic-based features are diverse and can be complementary. Based on these findings, we propose to utilize diverse semantic-based pretrained models to help each other for SVC. Besides, for joint usage of multiple models under mismatched time resolutions, we design the ReTrans strategy which owns high training efficiency and low inference latency. The experimental results verify the effectiveness of our proposed SVC framework based on Diverse Semantic-based Features Fusion, especially for the in-the-wild data of real-world applications.
In addition, this paper raises other questions worth exploring. Is it possible to pretrain one (or several) model(s) that can capture both melody and lyrics signals while being speaker-agnostic at the same time? How should we select the pretraining method and construct the pretraining corpus to improve the robustness? We will leave these to our future research.

### v3

@@ 3.3 Singing Voice Conversion Framework based on Diverse Semantic-based Features Fusion
During training, the DSFF-SVC conducts the reconstruction learning on the training corpus, and the speaker ID is just the speaker identity of every training sample. During conversion, given any source audio, we extract its semantic-based and energy features and stay them unchanged. To convert the speaker identity, we inject a reference speaker ID (which is seen in the training) to obtain the speaker-specific representations. For F0 features, we conduct the musical key transposition [^2] https://en.wikipedia.org/wiki/Transposition_(music) to make the reference speaker sing the source song in his vocal range. Specifically, following [31, 1], we shift the source F0 features by multiplying a factor, which is computed as the ratio between the F0 medians among the reference speaker’s training corpus and the source audio.
4 Experiments
We conduct experiments to answer the following questions:
  • EQ3: How effective is the resampling strategy compared to deep learning based method such as cross attention [21, 22] for fusing multiple features of mismatched time resolutions?
4.1 Experimental Setup
4.1.1 Evaluation Tasks
We adopt two conversion settings to evaluate: (1) Recording Studio Setting: following the most existing works, we utilize the high-quality singing corpus that is recorded in studio as the experimental data. The vocals are clean with virtually no noise or environmental interference. Specifically, we use Opencpop [32] as the target singer, whose training corpus is 5.2 hours of studio recorded singing voices. For source audios, we use M4Singer [33] and randomly 25 utterances for each timbre type respectively (including Soprano, Alto, Tenor, and Bass). (2) In-the-Wild Setting: in the real-world SVC application, usually the singing voices are separated from the background music, which will remain some artifacts or reverb in the vocals [11, 12]. We consider this as a more challenging conversion task to examine the robustness of the SVC systems. Specifically, we adopt a private corpus which contains 6.4 hours of singing voices of 15 professional singers (6 English singers and 9 Chinese singers). The vocals are separated by Ultimate Vocal Remover (UVR) [^3] https://github.com/Anjok07/ultimatevocalremovergui. We adopt four singers as the targets (an English male, an English female, a Chinese male, and a Chinese female). For source audios, we randomly sample 100 utterances which cover multiple musical genres including Pop, Rock, Folk and Soul.
4.1.2 Evaluation Metrics
For objective evaluation, following [1], we adopt Mel-cepstral distortion (MCD) [34], F0 Pearson correlation coefficient (F0CORR), F0 Root Mean Square Error (F0RMSE), Character Error Rate (CER) which is obtained with the recognition results of whisper-large ASR model [15]. Besides, following [35], we use the WavLM model [36] finetuned for speaker verification [^4] https://huggingface.co/microsoft/wavlm-base-plus-sv to compute the cosine speaker similarity score (SIM). For subjectivce evalution, we invite 12 volunteers who are experienced in the audio generation areas to conduct the Mean Opinion Score (MOS) evaluation in terms of naturalness and similarity. The naturalness score ranks from 1 (“Bad”) to 5 (“Excellent”), and the similarity score ranks from 1 (“Different speaker, sure”) to 4 (“Same speaker, sure”).
4.1.3 Base Models
  • TransformerSVC: It adopts a vanilla encoder-only transformer model[21] as the acoustic model. Its output is the mel-spectrogram.
  • VitsSVC: It is a VITS-based [37] model which is similar to the SoftVC-VITS [^5] https://github.com/svc-develop-team/so-vits-svc. It is an end-to-end framework and can directly produce waveform.
  • DiffWaveNetSVC: It adopts a diffusion-based acoustic model and could generate mel-spectrogram, which is proposed by [31]. The internal encoder of the diffusion framework is based on Bidirectional Non-Causal Dilated CNN[38], which is similar to WaveNet[39].
4.1.4 Implementation Details
For WeNet, we use the official models pretrained by 10k hours Wenetspeech [^6] https://github.com/wenet-e2e/wenet to extract semantic-based features. For Whisper, we use the multilingual MEDIUM model [^7] https://github.com/openai/whisper. For ContentVec, we use the official 500-CLASS model preatrained by 1k hours Librispeech [^8] https://github.com/auspicious3000/contentvec. The latent space dimension $D$ is set as 384. We adopt Parselmouth [40] to extract F0 and compute L2-norm of the amplitude of each short-time Fourier transform frame as energy features [41]. Following [31], we set their numbers of bins for quantization as 256. For the three base models, we adopt the implementations of Amphion [^9] https://github.com/open-mmlab/Amphion/tree/main/egs/svc [42]. For DiffWaveNetSVC and TransformerSVC, we use the pretrained Amphion Singing BigVGAN [^10] https://huggingface.co/amphion/BigVGAN_singing_bigdata as the vocoder to decode waveform from mel-spectrogram.
4.2 Performance of Different Semantic-based Features (EQ1)
In this section, we aim to explore the gap using semantic-based features alone to conduct the SVC task. We select DiffWaveNetSVC as the base model and conduct the task under the Recording Studio Setting. The experimental results are illustrated in Table 3.
@@ 4.2 Performance of Different Semantic-based Features (EQ1)
Fig. 3: The complementary role of diverse semantic-based features in melody modeling. More benefits of the joint usage of diverse semantic-based features (including spectrogram reconstruction, lyrics modeling, etc.) can be seen at our demo website.
On the one hand, the underlying knowledge of different semantic-based models can be distinct and diverse: (1) For modeling melody (F0CORR and F0RMSE), Whisper and ContentVec perform better than WeNet. This is mainly because the auxiliary tasks of the weak-supervised Whisper and the self-supervised ContentVec are beneficial for modeling more prosody-related signals. (2) For modeling lyrics (CER), Whisper performs best compared to the other two. It is assumed that the multilingual ASR tasks and the large-scale pretraining data allow Whisper to model more valid and robust semantic information. Moreover, ContentVec will provide better semantic signals than WeNet, although both are pretrained only on speech. It reveals that the self-supervised pretrained models could be more robust than the classical supervised ASR model. (3) When measuring speaker independence (SIM), the speaker similarity results of the three are comparable and all above 0.9, which is difficult to rank from human perception (see our demo page). This means that when using only semantic-based features for SVC, the three can all be considered speaker-agnostic.
@@ 4.3 Performance of the DSFF-SVC framework (EQ2)
+ Whisper 0.924 77.2 14.9% 0.881 0.848 183.8 18.7% 0.867
+ Whisper + ContentVec 0.931 75.5 16.2% 0.883 0.857 186.7 23.3% 0.868
@@
+ Whisper 0.945 144.4 17.8% 0.890 0.920 86.9 35.2% 0.869
+ Whisper + ContentVec 0.946 112.9 17.7% 0.886 0.921 79.5 32.3% 0.870
@@
+ Whisper 0.943 49.5 15.2% 0.884 0.921 73.6 21.1% 0.865
+ Whisper + ContentVec 0.940 55.2 15.7% 0.884 0.919 79.9 23.3% 0.867
Table 4: Objective Evaluation Results of the proposed DSFF-SVC framework.
To verify the generalization and robustness of the proposed DSFF-SVC framework, we conduct experiments based on both the recording studio and the in-the-wild settings for the three base models. Here we also use the auxiliary prosody features (including F0 and energy), to improve the melody modeling and the expressiveness of the system. The objective and subjective evaluation results can be seen in Table 4 and Table 5. It reveals that: (1) Conducting SVC under in-the-wild setting is more difficult. Facing to the more complex acoustic environment, only using the supervised WeNet produces a very poor performance. Integrating with the weak-supervised Whisper and the self-supervised ContentVec are usually beneficial, especially for Whisper. (2) For the generalization ability, we can observe that for all three models, the proposed idea of diversion semantic-based features fusion is effective in most cases under both settings. (3) For the robustness, particularly, under the in-the-wild setting, the improvement of integrating diverse semantic-based features is more obvious, especially for the objective CER and the subjective metrics.
Compared with Table 3 and 4, there are also many interesting observations about auxiliary prosody features (F0 and energy): (1) After injecting such prosody signals into SVC models, the intelligibility (CER) has been improved a lot. In our opinion, this is because the prosody in singing voice is expressive and characteristic, which is hard to be modeled without the explicit assistant of prosody features like F0. Therefore, using only semantic-based features, some tone- and intonation-related signals will not be learned well, leading to the worse intelligibility. (2) Except for the intelligibility, it is also notable about the change of conversion speaker similarity (SIM). After introducinh F0 features, the melody-related metrics have been improved a lot and the converted singing voices are not out of tune any more. In the meantime, however, the speaker similarity is also harmed. We suppose such auxiliary prosody features own some speaker-specific signals, e.g. the personalized vibrato patterns within F0 features. Therefore, directly copying the original prosody features from the source audio could not be the best choice for SVC, although it is a common way in the most SVC works.
Semantic-based Features Recording Studio In-the-Wild
Nat. ($\uparrow$) Sim. ($\uparrow$) Nat. ($\uparrow$) Sim. ($\uparrow$)
@@
+ Whisper 4.02 ±$\pm 0.18$ 3.13 ±$\pm 0.17$ 3.70 ±$\pm 0.18$ 2.86 ±$\pm 0.23$
+ Whisper + ContentVec 4.14 ±$\pm 0.19$ 3.25 ±$\pm 0.18$ 3.71 ±$\pm 0.18$ 2.82 ±$\pm 0.23$
Table 5: Subjective MOS evaluation results (with 95% confidence interval) of DiffWaveNetSVC. The full scores of Naturalness (Nat.) and Similarity (Sim.) are 5 and 4.
@@ 4.4 Performance of the Resolution Transformation based Features Fusion (EQ3)
RTX ($\uparrow$) RTF ($\downarrow$) F0CORR ($\uparrow$) CER ($\downarrow$) SIM ($\uparrow$)
Cross attention 191.9 2.57 0.896 16.4% 0.871
Resampling 415.0 0.86 0.936 15.8% 0.875
Table 6: The comparison between resampling and cross attention for features fusion.
To verify the performance and efficiency of the resampling for feature fusing, we select the cross attention [21] as the compared method. Specifically, we utilize only WeNet to extract semantic-based features and try different strategies to align its resolution to that of F0 and energy features. We conduct the experiments on DiffWaveNetSVC and follow NaturalSpeech2 [22] to adopt the cross attention for features fusion within the diffusion model.
Except for the quality evaluation for the conversion results, we also measure the computational cost of the both. Assume the duration of the whole training corpus is $d_{train}$, the duration of the evaluation samples to be converted is $d_{infer}$, the time to train an epoch is $t_{train}$, and the time to infer the evaluation samples is $t_{infer}$. We define RTX as $d_{train}/t_{train}$, which can measure the training efficiency. We define RTF as $t_{infer}/d_{infer}$ to measure the inference latency. The computational platform is a single NVIDIA Tesla V100.
The comparison results between resampling and cross attention are displayed in Table 6. We can see that when fusing the features of mismatched time resolutions for the SVC task, resampling is even slightly better than cross attention in terms of conversion quality. Moreover, as a non-learning method, resampling does not introduce any additional parameters and greatly accelerates both training efficiency (RTX) and inference speed (RTF) much.
@@ 5 Conclusion and Future Work
This paper investigates three semantic-based pretrained models of supervised, weak-supervised, and self-supervised fashions for SVC. We discover their capabilities including modeling melody and lyrics are diverse and can be complementary. Based on these findings, we propose the SVC framework based on Diverse Semantic-based Features Fusion. The experimental results confirm the effectiveness of our proposed framework, particularly for real-world applications involving in-the-wild data. In addition, this paper raises other questions worth exploring. Is it possible to pretrain one model that can capture both melody and lyrics signals while being speaker-agnostic at the same time? How should we select the pretraining method and construct the pretraining corpus to improve the robustness? We will leave these to our future research.